# Burgers equation for the bulk viscous pressure of quark matter

### v1

@@ I Introduction
In this Letter we focus on the bulk viscous effects of unpaired quark matter, that is, with the dissipation arising from compression and rarefaction. These effects play a crucial role in damping the rapid oscillations that follow the merger of compact stars [9, 10, 11], as well as various stellar oscillation modes in isolated stars [12, 13], both of which could influence gravitational wave signals [14]. Previous studies have computed bulk viscosity coefficients for different hadronic [15, 16, 17, 18, 19, 20] and quark matter phases [21, 22, 23, 24, 25] (see [26] for a review and a more complete set of references), typically as functions of the oscillation frequency, allowing quick estimates of energy dissipation. However, such formulations are not readily implementable in numerical simulations of the relativistic hydrodynamics associated with the relevant astrophysical settings. Only recently have efforts begun to incorporate bulk viscous dissipation into simulations of neutron star mergers, typically assuming that the bulk viscous stress tensor satisfies an Israel–Stewart (IS) equation [27, 28, 29, 30, 31, 32, 33, 34]. This second-order hydrodynamic framework introduces a relaxation time, an additional transport coefficient, in addition to the bulk viscosity itself, ensuring the causal propagation of hydrodynamical perturbations, in contrast to standard first-order formulations.
In this work we show that the bulk viscous pressure in unpaired quark matter evolves according to a Burgers-type equation, also used in viscoelastic media [35]. In [36] Gavassino has shown that the bulk pressure of systems with three independent chemical potentials and two distinct reaction rates, such as hadronic matter composed of neutrons, protons, electrons and muons, also follows Burgers-type evolution. While quark matter does not fulfill the above mentioned conditions, in this Letter we show that its bulk viscous pressure is also governed by a Burgers equation. Only in certain temperature and density regimes does it reduce to the more familiar IS form. We compute the full set of second-order transport coefficients entering the Burgers equation, conveniently expressed in terms of two partial bulk viscosity coefficients and two relaxation times. Although our formulation is general, we explicitly evaluate these transport coefficients for two distinct quark matter EOSs at finite densities, temperatures, and for a finite strange quark mass: one derived from perturbative QCD (pQCD), applicable at very high densities [37, 38, 39, 40, 41, 42], and another based on a modified MIT bag model [43, 44], more suitable for lower densities. Our approach provides a versatile framework that can accommodate more accurate EOS models or updated reaction rates as they become available, and can provide the proper description of the limiting IS case as well.
We use metric conventions (+,,,)(+,-,-,-)( + , - , - , - ) and natural units $k_{B}=\hbar=c=1$ throughout, unless otherwise stated.
@@ III Burgers equation
$\Pi$ $=$ $n_{u}\left(\delta\mu_{u}+\delta\mu_{e}-\delta\mu_{s}\right)+n_{d}\left(\delta\mu_{d}-\delta\mu_{s}\right)$ (5)
$+$ $n_{q}(\delta\mu_{s}-\frac{1}{3}\delta\mu_{e}),$
where $\delta\mu_{a}$ measures the deviation of the chemical potential of the particle $a$ with respect to its equilibrium value, and $n_{q}\equiv n_{u}+n_{d}+n_{s}$ is the quark density. Constraints of charge neutrality and conservation of baryon number have to be imposed (see Eqs. (27, 28) of End Matter). Consequently, the system is characterized by a single free chemical potential which we choose to be $\mu_{d}$.
@@ IV Results
where $G_{\mathrm{F}}\approx 1.166\times 10^{-5}\,\text{GeV}^{-2}$ is the Fermi coupling constant, $\Theta_{\mathrm{C}}\approx 13.02^{\circ}$ is the Cabibbo angle, $\alpha_{s}$ is the strong fine-structure constant, and $m_{s}$ is the strange quark mass, $m_{s}(2\,\mathrm{GeV})\approx 94\,\mathrm{MeV}$ [53]. Notably, the only rate scaling with a fifth power of the quark chemical potentials is $\lambda_{1}$, with the semileptonic processes remaining formally subdominant — nevertheless, we do not neglect them as they turn out to be crucial for understanding the bulk viscosity $\zeta_{-}$. We emphasize that in the expression for the rates the values for $m_{s}$ and $\alpha_{s}$ must be replaced by their model counterparts in the bag model. In the figures shown in the main text, we have always fixed $m_{s}$ to its bare value and set the bag model coupling $a_{4}\approx 1-2\alpha_{s}/\pi$ [54] to $a_{4}=0.7$. We discuss the parameter dependence in more detail in the End Matter.
@@
The bulk viscosities $\zeta_{\pm}$ behave similarly (see Fig. 2), but $\zeta_{-}$ saturates to a nearly constant value, and starting at approximately $T=0.1~{}\,\mathrm{MeV}$, $\zeta_{+}$ overtakes $\zeta_{-}$. The pQCD error bands, obtained by varying the renormalization scale (see the End Matter for details), are sizable for $\zeta_{-}$ before the inflection point, and considerably more modest for all other quantities.
@@ V Discussion
Finally, let us comment that while the Burgers equation (14) describes the close-to-chemical-equilibrium evolution of quark matter, the same effects could be described by taking into account the individual evolution of the densities of the quarks and electrons, after imposing the constraints of charge neutrality and baryon number conservation in their evolution. However, our proposed formulation is considerably simpler and will be of great use in analytical models and could facilitate numerical studies of both stellar oscillation modes and mergers of quark or hybrid stars.
@@ VI Acknowledgements
VII End Matter
VII.1 Frequency-dependent bulk viscosity
@@
$\kappa_{3}=1/(\tau_{+}\tau_{-})^{2},$ (22)
$\kappa_{1}/\kappa_{3}=\zeta_{+}+\zeta_{-},\qquad\kappa_{4}/\kappa_{3}=\tau_{+}^{2}+\tau^{2}_{-},$ (23)
$\kappa_{2}/\kappa_{3}=(\tau_{+}+\tau_{-})(\zeta_{+}\tau_{-}+\zeta_{-}\tau_{+}).$ (24)
We have checked that with the values of the second-order coefficients found in this manuscript we reproduce the frequency-dependent bulk viscosity computed in [25] [^3] Note that in [25], the chemical potential $\mu_{2}$ is denoted $-\mu_{3}$, and the chemical potential $\mu_{2}$ found there corresponds to our linear combination $\mu_{2}-\mu_{1}$. The rates $\lambda_{2,3}$ are likewise swapped.. One can even use the above relations to obtain the associated Burgers equation fulfilled by any system that has a frequency-dependent bulk viscosity with the same dependence as in Eq. (21).
@@
VII.2 Details of constraints and the equations of state
@@
00 0 $=\delta n_{Q}=\sum_{a}\left(\chi_{e}^{\ a}-\chi_{u}^{\ a}\right)\delta\mu_{a},$ (27)
00 0 $=\delta n_{q}=\sum_{a}\left(\chi_{d}^{\ a}+\chi_{s}^{\ a}-\chi_{e}^{\ a}\right)\delta\mu_{a}.$ (28)
@@
[…]
Here, the $m_{s}$-parameter should be thought to represent the constituent quark mass, and acts as a third model parameter. Indeed, as the parameter responsible for breaking conformal invariance, it is the one most relevant for bulk viscosities. Densities and therefore also susceptibilities are independent of the bag constant, whereas the pressure depends approximately linearly on $a_{4}$. For small variations of $a_{4}$, its impact is essentially imperceptible on the doubly logarithmic scale used here for the bulk viscosities. To be more quantitative, we set $a_{4}=0.7$ [14] having checked explicitly that varying it between $0.6-0.8$ changes $\tau_{+}$ by $\lesssim 15\%$, $\zeta_{+}$ by $\lesssim 30\%$, and both $\tau_{-}$ and $\zeta_{-}$ orders of magnitude less.
@@
Appendix A SUPPLEMENTAL MATERIAL
@@
$-\left(\chi^{-1}\right)_{d}^{\ d}\left(\chi_{A}^{-1}-\chi_{B}^{-1}-\chi_{C}^{-1}\right)-2\left(\chi^{-1}\right)_{s}^{\ s}\chi_{B}^{-1}+\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\chi_{B}^{-1}\bigg{\}}$
@@
$+\left(\chi^{-1}\right)_{d}^{\ d}\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)+\left(\chi^{-1}\right)_{s}^{\ s}\left(\chi_{A}^{-1}+\chi_{B}^{-1}-\chi_{C}^{-1}\right)-\left(\chi_{A}^{-1}-\chi_{B}^{-1}-\chi_{C}^{-1}\right)\chi_{A}^{-1}\bigg{\}}$
@@
+$+\Bigg{\{}-\left(\chi_{A}^{-1}-\chi_{B}^{-1}\right)^{2}+\left(\chi_{A}^{-1}+\chi_{B}^{-1}\right)\chi_{C}^{-1}-\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\left[\left(\chi_{u}^{\ u}\right)^{-1}+\frac{1}{3}\left(\chi_{e}^{\ e}\right)^{-1}\right]$
@@
+$+\Bigg{\{}\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\chi_{B}^{-1}+\left(\chi_{A}^{-1}+\chi_{B}^{-1}-\chi_{C}^{-1}\right)\left[\left(\chi_{u}^{\ u}\right)^{-1}+\frac{1}{3}\left(\chi_{e}^{\ e}\right)^{-1}\right]$
@@
+$+\Bigg{\{}-\left(\chi_{A}^{-1}-\chi_{B}^{-1}-\chi_{C}^{-1}\right)\chi_{A}^{-1}-2\chi_{A}^{-1}\left[\left(\chi_{u}^{\ u}\right)^{-1}+\frac{1}{3}\left(\chi_{e}^{\ e}\right)^{-1}\right]$
$+\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\left(\chi_{d}^{\ d}\right)^{-1}+\left(\chi_{A}^{-1}+\chi_{B}^{-1}-\chi_{C}^{-1}\right)\left(\chi_{s}^{\ s}\right)^{-1}\Bigg{\}}n_{q}\lambda_{1}\Pi_{2},$
@@
+$+\left\{n_{q}\left[\left(\chi^{-1}\right)_{d}^{\ d}-\left(\chi^{-1}\right)_{u}^{\ u}-\frac{1}{3}\left(\chi^{-1}\right)_{e}^{\ e}\right]+\left(n_{d}-n_{u}\right)\chi_{B}^{-1}-n_{s}\left(\chi_{A}^{-1}-\chi_{C}^{-1}\right)\right\}\Pi_{2},$ (33)

### v2

@@ I Introduction
In this Article we focus on the bulk viscous effects of unpaired quark matter, that is, with the dissipation arising from compression and rarefaction. These effects play a crucial role in damping the rapid oscillations that follow the merger of compact stars [9, 10, 11], as well as various stellar oscillation modes in isolated stars [12, 13], both of which could influence gravitational wave signals [14]. Previous studies have computed bulk viscosity coefficients for different hadronic [15, 16, 17, 18, 19, 20] and quark matter phases [21, 22, 23, 24, 25] (see [26] for a review and a more complete set of references), typically as functions of the oscillation frequency, allowing quick estimates of energy dissipation. However, such formulations are not readily implementable in numerical simulations of the relativistic hydrodynamics associated with the relevant astrophysical settings. Only recently have efforts begun to incorporate bulk viscous dissipation into simulations of neutron star mergers, typically assuming that the bulk viscous stress tensor satisfies an Israel–Stewart (IS) equation [27, 28, 29, 30, 31, 32, 33, 34]. This second-order hydrodynamic framework introduces a relaxation time, an additional transport coefficient, in addition to the bulk viscosity itself, ensuring the causal propagation of hydrodynamical perturbations, in contrast to standard first-order formulations.
In this work we show that the bulk viscous pressure in unpaired quark matter evolves according to a Burgers-type equation, also used in viscoelastic media [35]. In [36] Gavassino has shown that the bulk pressure of systems with three independent chemical potentials and two distinct reaction rates, such as hadronic matter composed of neutrons, protons, electrons and muons, also follows Burgers-type evolution. While quark matter does not fulfill the above mentioned conditions, in this Article we show that its bulk viscous pressure is also governed by a Burgers equation. Only in certain temperature and density regimes does it reduce to the more familiar IS form. We compute the full set of second-order transport coefficients entering the Burgers equation, conveniently expressed in terms of two partial bulk viscosity coefficients and two relaxation times. Although our formulation is general, we explicitly evaluate these transport coefficients for two distinct quark matter EOSs at finite densities, temperatures, and for a finite strange quark mass: one derived from perturbative QCD (pQCD), applicable at very high densities [37, 38, 39, 40, 41, 42], and another based on a modified MIT bag model [43, 44], more suitable for lower densities. Our approach provides a versatile framework that can accommodate more accurate EOS models or updated reaction rates as they become available, and can provide the proper description of the limiting IS case as well.
We use metric conventions (+,,,)(+,-,-,-) and natural units $k_{B}=\hbar=c=1$ throughout, unless otherwise stated.
@@ III Burgers equation
$\Pi$ $=n_{u}\left(\delta\mu_{u}+\delta\mu_{e}-\delta\mu_{s}\right)$
$+n_{d}\left(\delta\mu_{d}-\delta\mu_{s}\right)+n_{q}(\delta\mu_{s}-\frac{1}{3}\delta\mu_{e}),$ (5)
where $\delta\mu_{a}$ measures the deviation of the chemical potential of the particle $a$ with respect to its equilibrium value, and $n_{q}\equiv n_{u}+n_{d}+n_{s}$ is the quark density. Constraints of charge neutrality and conservation of baryon number have to be imposed (see Eqs. (25, 26) of Appendix B) to express $\Pi$ in terms of the algebraically independent chemical potentials $\mu_{1},\mu_{2}$. This amounts to writing the chemical potentials appearing in Eq. (5) in terms of $\mu_{1},\mu_{2}$, densities $n_{a}$, and (inverse) susceptibilities $\chi_{ab}^{(-1)}$. As a consequence of the constraints, the equilibrium system is characterized by a single free chemical potential which we choose to be $\mu_{d}$.
@@ IV Results
where $G_{\mathrm{F}}\approx 1.166\times 10^{-5}\,\text{GeV}^{-2}$ is the Fermi coupling constant, $\Theta_{\mathrm{C}}\approx 13.02^{\circ}$ is the Cabibbo angle, $\alpha_{s}$ is the strong fine-structure constant, and $m_{s}$ is the strange quark mass, $m_{s}\left(\Lambda=2\,\mathrm{GeV}\right)\approx 94\,\mathrm{MeV}$ [53], evaluated at a fixed renormalization scale $\Lambda$ and evolved according to the renormalization group equations. Notably, the only rate scaling with a fifth power of the quark chemical potentials is $\lambda_{1}$, with the semileptonic processes remaining formally subdominant — nevertheless, we do not neglect them as they turn out to be crucial for understanding the bulk viscosity $\zeta_{-}$. We emphasize that in the expression for the rates the values for $m_{s}$ and $\alpha_{s}$ must be replaced by their model counterparts in the bag model. In the figures shown in the main text, we have always fixed $m_{s}$ to its bare value and set the bag model coupling $a_{4}\approx 1-2\alpha_{s}/\pi$ [54] to $a_{4}=0.7$. We discuss the parameter dependence in more detail in Appendix B.
@@
The bulk viscosities $\zeta_{\pm}$ behave similarly (see Fig. 2), but $\zeta_{-}$ saturates to a nearly constant value, and starting at approximately $T=0.1~\,\mathrm{MeV}$, $\zeta_{+}$ overtakes $\zeta_{-}$. The pQCD error bands, obtained by varying the renormalization scale (see Appendix B for details), are sizable for $\zeta_{-}$ before the inflection point, and considerably more modest for all other quantities.
@@ V Discussion
Finally, let us comment that while the Burgers equation (14) describes the close-to-chemical-equilibrium evolution of quark matter, the same effects could be described by taking into account the individual evolution of the densities of the quarks and electrons, after imposing the constraints of charge neutrality and baryon number conservation in their evolution. However, our proposed formulation is considerably simpler and will be of great use in analytical models and could facilitate numerical studies of both stellar oscillation modes and mergers of quark or hybrid stars. In future, the treatment found here could also be further extended by considering the variety of superconducting phases found in dense matter.
@@ VI Acknowledgements
VII Data Availability
The data that support the findings of this article are openly available [60].
Appendix A Frequency-dependent bulk viscosity
@@
$\kappa_{3}$ $=\frac{1}{(\tau_{+}\tau_{-})^{2}},\qquad\frac{\kappa_{1}}{\kappa_{3}}=\zeta_{+}+\zeta_{-},$
$\qquad\frac{\kappa_{4}}{\kappa_{3}}$ $=\tau_{+}^{2}+\tau^{2}_{-},\qquad\,\,\frac{\kappa_{2}}{\kappa_{3}}=\zeta_{+}\tau_{-}^{2}+\zeta_{-}\tau_{+}^{2}.$ (22)
We have checked that with the values of the second-order coefficients found in this manuscript we reproduce the frequency-dependent bulk viscosity computed in [25] [^3] Note that in [25], the chemical potential $\mu_{2}$ is denoted $\mu_{3}$, and the chemical potential $\mu_{2}$ found there corresponds to our linear combination $\mu_{2}-\mu_{1}$. The rates $\lambda_{2,3}$ are likewise swapped.. One can even use the above relations to obtain the associated Burgers equation fulfilled by any system that has a frequency-dependent bulk viscosity with the same dependence as in Eq. (21).
@@
Appendix B Details of constraints and the equations of state
@@
$0$ $=\delta n_{Q}=\sum_{a}\left(\chi_{e}^{\ a}-\chi_{u}^{\ a}\right)\delta\mu_{a},$ (25)
$0$ $=\delta n_{q}=\sum_{a}\left(\chi_{d}^{\ a}+\chi_{s}^{\ a}-\chi_{e}^{\ a}\right)\delta\mu_{a}.$ (26)
@@
$p_{s}=\frac{3T}{\pi^{2}}\int_{0}^{\infty}\mathrm{d}kk^{2}\,\Bigg\{\ln\left[1+\exp\left(-\frac{\sqrt{k^{2}+m_{s}^{2}}-\mu_{s}}{T}\right)\right]+\ln\left[1+\exp\left(-\frac{\sqrt{k^{2}+m_{s}^{2}}+\mu_{s}}{T}\right)\right]\Bigg\}.$ (28)
Here, the $m_{s}$-parameter should be thought to represent the constituent quark mass, and acts as a third model parameter. Indeed, as the parameter responsible for breaking conformal invariance, it is the one most relevant for bulk viscosities. Densities and therefore also susceptibilities are independent of the bag constant, whereas the pressure depends approximately linearly on $a_{4}$. For small variations of $a_{4}$, its impact is essentially imperceptible on the doubly logarithmic scale used here for the bulk viscosities. To be more quantitative, we set $a_{4}=0.7$ [14] having checked explicitly that varying it between 0.60.80.6-0.8 changes $\tau_{+}$ by $\lesssim 15\%$, $\zeta_{+}$ by $\lesssim 30\%$, and both $\tau_{-}$ and $\zeta_{-}$ orders of magnitude less.
@@
Appendix C Explicit formulae
@@
$-\left(\chi^{-1}\right)_{d}^{\ d}\left(\chi_{A}^{-1}-\chi_{B}^{-1}-\chi_{C}^{-1}\right)-2\left(\chi^{-1}\right)_{s}^{\ s}\chi_{B}^{-1}+\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\chi_{B}^{-1}\bigg\}$ (30)
@@
$+\left(\chi^{-1}\right)_{d}^{\ d}\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)+\left(\chi^{-1}\right)_{s}^{\ s}\left(\chi_{A}^{-1}+\chi_{B}^{-1}-\chi_{C}^{-1}\right)-\left(\chi_{A}^{-1}-\chi_{B}^{-1}-\chi_{C}^{-1}\right)\chi_{A}^{-1}\bigg\}$ (31)
@@
$+\Bigg\{-\left(\chi_{A}^{-1}-\chi_{B}^{-1}\right)^{2}+\left(\chi_{A}^{-1}+\chi_{B}^{-1}\right)\chi_{C}^{-1}-\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\left[\left(\chi_{u}^{\ u}\right)^{-1}+\frac{1}{3}\left(\chi_{e}^{\ e}\right)^{-1}\right]$
@@
$+\Bigg\{\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\chi_{B}^{-1}+\left(\chi_{A}^{-1}+\chi_{B}^{-1}-\chi_{C}^{-1}\right)\left[\left(\chi_{u}^{\ u}\right)^{-1}+\frac{1}{3}\left(\chi_{e}^{\ e}\right)^{-1}\right]$
@@
$+\Bigg\{-\left(\chi_{A}^{-1}-\chi_{B}^{-1}-\chi_{C}^{-1}\right)\chi_{A}^{-1}-2\chi_{A}^{-1}\left[\left(\chi_{u}^{\ u}\right)^{-1}+\frac{1}{3}\left(\chi_{e}^{\ e}\right)^{-1}\right]$
$+\left(\chi_{A}^{-1}-\chi_{B}^{-1}+\chi_{C}^{-1}\right)\left(\chi_{d}^{\ d}\right)^{-1}+\left(\chi_{A}^{-1}+\chi_{B}^{-1}-\chi_{C}^{-1}\right)\left(\chi_{s}^{\ s}\right)^{-1}\Bigg\}n_{q}\lambda_{1}\Pi_{2},$ (33)
@@
$+\left\{n_{q}\left[\left(\chi^{-1}\right)_{d}^{\ d}-\left(\chi^{-1}\right)_{u}^{\ u}-\frac{1}{3}\left(\chi^{-1}\right)_{e}^{\ e}\right]+\left(n_{d}-n_{u}\right)\chi_{B}^{-1}-n_{s}\left(\chi_{A}^{-1}-\chi_{C}^{-1}\right)\right\}\Pi_{2},$ (34)